Pushing over the file cabinet to keep out the

zombies requires 1000 N of force. The cabinet

accelerates at 20 m/s

2

. What is the mass of the

cabinet?


1
Expert's answer
2022-02-18T15:53:11-0500

From Newton's second law "F=m \cdot a"

Let's write down

"m=\frac{F}{a}=\frac{1000}{20}=50kg"